Assume the probability is 1/2 that a child born at any time is a boy and births are independent. What is the probability of the event that if a family has 5 children (no twins), none of them are boys. I am stressing out here. I think I am overthinking these word problems.
Answer by greenestamps(13327)



No boys means all girls; P(boy)=1/2 means P(girl)=1/2.

P(5 girls)=(1/2)^5 = 1/32.